INFRARED SPECTRA OF OCS-C_{6}H_{6}, OCS-C_{6}H_{6}-HE AND OCS-C_{6}H_{6}-NE VAN DER WAALS COMPLEXES

M. DEHGHANY, \underline{J. NOROOZ OLIAEE}, MAHIN AFSHARI, N. MOAZZEN-AHMADI, Department of Physics and Astronomy, University of Calgary, 2500 University Dr., N.W., Calgary, Alberta T2N 1N4, Canada; A.R.W. McKELLAR, Steacie Institute for Molecular Sciences, National Research Council of Canada, Ottawa, Ontario, K1A 0R6, Canada.

The infrared spectrum of weakly-bound OCS-C_{6}H_{6} is studied in the region of the ν_{1} fundamental band of OCS (∼2060 cm^{−1}) using a tunable diode laser spectrometer to probe a pulsed supersonic jet expansion. A very simple band is observed, corresponding to a parallel transition of a symmetric top. It is shifted by -11.1 cm^{−1} with respect to the free OCS monomer. The resulting structure has OCS located along the benzene C_{6} symmetry axis in an S-bonded configuration with a center of mass separation of 4.42 Å, in good agreement with previous microwave spectra.^{a} The isotopomers OCS-^{13}C^{12}C_{5}H_{6} and OC^{34}S-C_{6}H_{6} are also observed. Similar bands are observed for the OCS-C_{6}H_{6}-He and OCS-C_{6}H_{6}-Ne trimers, whose structure is obtained by adding an on-axis rare gas atom to the other side of the benzene. However, the analogous band for OCS-C_{6}H_{6}-Ar was not detected, raising the possibility that the stable form of this trimer may not have the same symmetrical structure.

aU. Dahmen, H. Dreizler, and W. Stahl, Ber. Bunsenges. Phys. Chem. 99, 434 (1995).